OpenStudy (anonymous):

What is the equation of the trend line? (Round the slope to the nearest hundredth.) A. y = 1.25x B. y = 15.00x C. y = 0.80x D. y = 12.00x

OpenStudy (igreen):

Oh, that's (25, 20)..my bad.

OpenStudy (igreen):

The line passes through the points (15, 12) and (25, 2-) Plug them into the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{20-12}{25-15}\) Simplify

OpenStudy (igreen):

Can you simplify that?

OpenStudy (anonymous):

ya thanks....